OpenStudy (mef4.0):

Evaluate: lim (x^2 +6x +5)/(x^2-3x-4) x--> -1

OpenStudy (anonymous):

Okay so here you have to factor the function so that you can cancel out some parts of the equation and avoid dividing by zero.

OpenStudy (anonymous):

LagrangeSon, you cannot, because in that case you would be dividing by 0. Rather, factor the two polynomials:

OpenStudy (anonymous):

lim x-> =-1 (x+1)(x+5)/(x-4)(x+1)

OpenStudy (anonymous):

Now you can cancel out the (x+1)'s in the numerator and denominator and plug in -1 to get your answer.

OpenStudy (anonymous):

And, you should get, 4/-5 = -4/5

OpenStudy (mef4.0):

so - is the answer---f(x) = - 4/5 as x approaches -1? this sounds stupid, but I'm not sure how to notate it- what's the correct way to write the answer?

OpenStudy (anonymous):

It's not a stupid question -- in most Calculus classes professors will penalize you harshly for incorrect notation. Yes, I would say that: lim write the function here = (-1+5)/(-1-4)= 4/-5 = -4/5 x->-1

OpenStudy (anonymous):

The answer is just a number.

OpenStudy (mef4.0):

thanks very much

OpenStudy (anonymous):

You're welcome. Good luck on your final. (I have a Calc I final tomorrow, haha).

OpenStudy (mef4.0):

Where is your final? - what school?